# LayLens: Improving Deepfake Understanding through Simplified Explanations

### v1

@@ Abstract.
Deepfake Detection, Explanation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; July 2025; Melbourne, Australiabooktitle: ICMI ’25: INTERNATIONAL CONFERENCE ON MULTIMODAL INTERACTION, October 13–17, 2025, Canberra, Australiasubmissionid: 1003ccs: Human-centered computing Graphical user interfaces
@@ 2. Related Work
Traditional deepfake detection models, such as  (Zhou et al., 2018; Matern et al., 2019; Chugh et al., 2020) primarily focused on obtaining high accuracy in detecting manipulated inputs but lacked intuitive interpretability for lay users. In earlier years, explainability in deepfake detection was introduced through saliency-based techniques like LRP (Bach et al., 2015), Grad-CAM (Selvaraju et al., 2017), LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017), which produce saliency maps highlighting image regions most influential to the classifier’s output. These approaches are model-agnostic and were adopted to identify the important input regions in deepfake detection models in works such as  (Pino et al., 2021; Malolan et al., 2020). Later works such as  (Soltandoost et al., 2025) propose identifying embeddings corresponding to local facial regions in images, and use them to produce interpretable classification outputs. Recently, Vision-Language model based approaches, such as FakeShield  (Xu et al., 2025) and SIDA  (Huang et al., 2025) have become popular, which provide textual rationale along with localization of the tampered region. Despite these advancements, current explainable detection tools exhibit gaps that limit their usefulness for non-experts. A common issue is that the explanations are too low-level or technical. There is an evident gap in the literature for a deepfake detection interface that provides simplified yet informative visual and textual explanations accessible to the general public.
@@ 3. Interface Design
LayLens is designed as an interactive, end-to-end system that not only detects deepfake images but also presents explanations in a form accessible to both technical and non-technical users. The system integrates state-of-the-art methods in deepfake detection, localization and explanation, along with generative image editing, and wraps them in a user-friendly interface that encourages interpretability and engagement. After the user uploads an image to the interface, the following components are triggered:
@@ 3.0.2. Explanation Card:
    Simplified Explanation (Figure 3): Translates the technical rationale into natural language, tailored for non-expert users. Each identified region (e.g., “Central Teddy Bear”) is paired with a relevant emoji and an accessible explanation. Specifically, we employ Qwen-VL (Bai et al., 2025), a foundation Vision Language Model. It is prompted with both the image and FakeShield’s technical explanation and tasked with generating a structured, human-readable JSON output. For each manipulated region, Qwen-VL outputs:
@@
Thus, LayLens offers a complete workflow that adapts to users’ preferred level of technical detail while always grounding its decisions in visual evidence. It enhances the accessibility, transparency, and interpretability of deepfake detection for a broad spectrum of users, from forensic analysts to everyday citizens. This design lets different users choose their preferred explanation detail level, while always seeing a visual demonstration of the AI’s reasoning.
@@ 4. User Survey
[…]
To evaluate the effectiveness of LayLens in providing accessible explanations for deepfake images, we conducted a user study with 15 participants, of whom 11 were familiar with the concept of deepfakes, and 3 had prior experience with deepfake detection tools. Each participant interacted with the system by analyzing 10 AI-manipulated images. Overall, users preferred the simplified explanations over the complex ones in 65.3% of the cases. Notably, in 81.3% of comparisons, participants reported that the simplified explanations reduced their cognitive load in understanding why an image might be a deepfake. The side-by-side visualization, featuring the uploaded (potentially fake) image alongside a plausibly reconstructed original, was considered helpful in enhancing understanding in 69.3% of the instances. Furthermore, 80% of participants indicated that the experience improved their confidence in detecting deepfakes in the future, and 93.3% expressed interest in using such a tool for identifying manipulated media going forward. The distribution of responses across all questions, measured using a 5-point Likert scale, is illustrated in Figure 4. Additionally, we performed a Wilcoxon signed-rank test to assess perceived changes in Ease of Understanding, Clarity and Accuracy when switching from complex to simplified explanations. The resulting p-values were $063.25e-$ 06, $0.01$ and $0.30$, respectively. These results indicate statistically significant improvements (p $\leq 0.05$) in both Ease of Understanding and Clarity when simplified explanations are presented.
@@ 5. Conclusion
In this work, we presented LayLens, a system to make deepfake detection more accessible, interpretable and engaging for users with varying levels of technical expertise. By integrating detection, natural language simplification, and generative reconstruction, LayLens enables both technical and non-expert users to understand why an image may be fake and what the original might have looked like. Our user study shows that simplified, visually grounded explanations reduce cognitive load and enhance user confidence. LayLens demonstrates that deepfake detection can be made both accurate and accessible, paving the way for more transparent and trustworthy media forensics. In future, we will add performance comparisons of this tool using proprietary AI tools such as ChatGPT (OpenAI, 2024) and Gemini (Anil and et al., 2023); and work towards creating an open-source, efficient VLM to directly obtain simplified explanations.

### v2

@@ Abstract.
Deepfake Detection; Explanation
$c$s: Human-centered computing Graphical user interfacesccs: Computing methodologies Biometricsccs: Human-centered computing Accessibility systems and toolsccs: Social and professional topics Assistive technologies
@@ 2. Related Work
Traditional deepfake detection models, such as  (Zhou et al., 2018; Matern et al., 2019; Chugh et al., 2020) primarily focused on obtaining high accuracy in detecting manipulated inputs but lacked intuitive interpretability for lay users. In earlier years, explainability in deepfake detection was introduced through saliency-based techniques like LRP (Bach et al., 2015), Grad-CAM (Selvaraju et al., 2017), LIME (Ribeiro et al., 2016) and SHAP (Lundberg et al., 2017), which produce saliency maps highlighting image regions most influential to the classifier’s output. These approaches are model-agnostic and were adopted to identify the important input regions in deepfake detection models in works such as  (Pino et al., 2021; Malolan et al., 2020). Later works such as  (Soltandoost et al., 2025) propose identifying embeddings corresponding to local facial regions in images, and use them to produce interpretable classification outputs. Recently, Vision-Language model based approaches, such as FakeShield  (Xu et al., 2025) and SIDA  (Huang et al., 2025) have become popular, which provide textual rationale along with localization of the tampered region. Despite these advancements, current explainable detection tools exhibit gaps that limit their usefulness for non-experts. A common issue is that the explanations are too low-level or technical. There is an evident gap in the literature for a deepfake detection interface that provides simplified yet informative visual and textual explanations accessible to the general public.
@@ 3. Interface Design
[…]
LayLens is designed as an interactive, end-to-end system that not only detects deepfake images but also presents explanations in a form accessible to both technical and non-technical users. The system integrates state-of-the-art methods in deepfake detection, localization and explanation, along with generative image editing, and wraps them in a user-friendly interface that encourages interpretability and engagement. (Figure 1) gives an overview of this three-stage pipeline, showing how LayLens moves from deepfake detection, through explanation simplification, to visual reconstruction. After the user uploads an image to the interface, the following components are triggered:
@@ 3.0.2. Explanation Card:
    Simplified Explanation (Figure 4): Performs automatic text simplification on complex explanation. Each identified region (e.g., “Central Teddy Bear”) is paired with a relevant emoji and a simplified explanation, generated by a VLM (Bai et al., 2025). It is prompted with both the image and FakeShield’s technical explanation and tasked with generating a structured, human-readable JSON output. For each manipulated region, the VLM outputs:
@@
Thus, LayLens offers a complete workflow that lets the users choose their preferred level of technical detail while always grounding its decisions in visual evidence. It enhances the accessibility, transparency, and interpretability of deepfake detection for a broad spectrum of users, from forensic analysts to everyday citizens. This design lets different users choose their preferred explanation detail level, while always seeing a visual demonstration of the AI’s reasoning.
@@ 4. User Survey
To evaluate the effectiveness of LayLens in providing accessible explanations for deepfake images, we conducted a user study with 15 participants, of whom 11 were familiar with the concept of deepfakes, and 3 had prior experience with deepfake detection tools. Each participant interacted with the system by analyzing 10 AI-manipulated images. Overall, users preferred the simplified explanations over the complex ones in 65.3% of the cases. Notably, in 81.3% of comparisons, participants reported that the simplified explanations reduced their cognitive load in understanding why an image might be a deepfake. The side-by-side visualization, featuring the uploaded (potentially fake) image alongside a plausibly reconstructed original, was considered helpful in enhancing understanding in 69.3% of the instances. Furthermore, 80% of participants indicated that the experience improved their confidence in detecting deepfakes in the future, and 93.3% expressed interest in using such a tool for identifying manipulated media going forward. The distribution of responses across all questions, measured using a 5-point Likert scale, is illustrated in Figure 5. Additionally, we performed a Wilcoxon signed-rank test to assess perceived changes in Ease of Understanding, Clarity and Accuracy when switching from complex to simplified explanations. The resulting p-values were 3.25e063.25e-06, $0.01$ and $0.30$, respectively. These results indicate statistically significant improvements (p $\leq 0.05$) in both Ease of Understanding and Clarity when simplified explanations are presented.
@@ 5. Conclusion
In this work, we presented LayLens, a system to make deepfake detection more accessible, interpretable and engaging for users with varying levels of technical expertise. By integrating detection, natural language simplification, and generative reconstruction, LayLens enables both technical and non-expert users to understand why an image may be fake and what the original might have looked like. Our user study shows that simplified, visually grounded explanations reduce cognitive load and enhance user confidence. LayLens demonstrates that deepfake detection can be made both accurate and accessible, paving the way for more transparent and trustworthy media forensics. As a future work, we will develop an open-source, efficient and explainable model for large-scale audio-visual deepfakes (Cai et al., 2025) and further, extend it to incorporate multi-lingual, code-switched videos (Kuckreja et al., 2025).